What pressure (in atm) will 0.52 moles of N2 exert in a 2.6 L container at 25°C?
Answer #1

The equation to be used to solve the equations is the ideal gas equation:

PV = nRT where P = pressure(in atm), V = volume(in L) , n = number of moles of gas, R = Universal Gas Constant ( 0.0821 L atm mol-1 K-1) and T = Temperature (in K).

Here,

n = 0.52 moles

V = 2.6 L

T = 25 C = ( 25 + 273 ) K = 298 K

Putting the values:

P (2.6 L) = (0.52 moles) (0.0821 L atm mol-1 K-1) (298 K)

Solving, P = 4.9 atm ( answer)